OpenStudy (anonymous):

You saved $20,000.00 and want to diversify your monies. You invest 45% in a Treasury bond for 3 years at 4.35% APR compounded annually. You place 15% in a CD at 3.75% APR for 3 years compounded annually. 20% you invest in a stock plan and the remainder is in a savings account at 2.90% APR compounded annually. The stock plan increases 8% the first year, decreases in value by 4% the second year, and increases by 6% the third year.

OpenStudy (anonymous):

1. are the balances for each type of investment at the end of the third year? 2. What is your total gain from all of the investments combined? 3. If you had invested 45% in stock What and 20% in Treasury bonds, would you have more or less of a gain after the three years?

OpenStudy (anonymous):

First you need to calculate how much money goes to each type 20000*0.45 = 9000 in the bond 20000*0.15 = 3000 in the CD 20000*0.20 = 4000 in stocks 20000*0.029 = 580 in savings Then for each type use the given data and the formula A = P(1 + r/n)^(nt) to find the total. For example for the bonds your P is $9000, r is 0.0435, n = 1, and t = 3 So the formula becomes \[A = 9000*(1+\frac{ 0.0435 }{ 1 })^{3}\ = $10,266.33] and the rest are the same with different numbers

OpenStudy (anonymous):

so i need a little more help because that was somewhat confusing to me so do i use that problem on all of the 3 questions or just the first one? @LastTrainHome22

OpenStudy (anonymous):

That equation will work for all 4 cases just by changing the numbers you use. For the Bond, the CD, and the savings account all you need to do is change the P to be whatever amount of money you give to each account, and r becomes whatever interest rate is specified in the problem statement

OpenStudy (anonymous):

For the stocks account, however, you need to calculate it 3 times since the interest rate changes each year. In this case you would calculate the first year using P=4000, r is the first year interest rate, so thats 0.08, and the time t is 1. For the second year you would set P equal to whatever you calculated for the end of year one, and change r to the new interest rate of -0.04. Then for the third year change P to whatever you just calculated and change r to the third year interest rate of 0.06

OpenStudy (anonymous):

sorry one more question what would n be @LastTrainHome22

OpenStudy (anonymous):

I got 4396.03 but you may want to check my work

OpenStudy (anonymous):

thank u so much u were so helpful